# MHE under parametric uncertainty - Robust state estimation without informative data

### v1

@@ 5 Discussion
In this section, we discuss the relation of the MHE approach presented in Section 4 to various existing methods for state and parameter estimation from the literature.
@@ 5.1 Standard Moving Horizon Estimation
Standard MHE approaches for state and parameter estimation similar to the one described in Section 2.3 are widely applied [46, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Assuming PE, robust convergence of the joint state and parameter estimation error system under a standard MHE scheme can be established [47, Thm. 2]. However, as seen in Section 3, lack of PE can lead to parameter drift and unstable state estimates. We overcome this issue by replacing the standard prior weighting on the parameter with a regularization with respect to a (constant) prior parameter estimate $\bar{\theta}_{0}$. This ensures practical stability of the state estimates irrespective of PE. In case of sufficiently large PE, the standard MHE formulation ensures robust stability, while the proposed MHE may result in less accurate state estimates due to the bias with respect to the prior parameter. Hence, while the proposed modification improves stability properties independent of PE, it may also deteriorate the estimation performance if PE is sufficiently large. Deriving an MHE formulation that combines both properties remains an interesting question for future research.
@@ 5.3 Adaptive Estimation
Classically, adaptive observers are applied for state estimation in case of parametric uncertainty. While there exist adaptive observers with guarantees in the absence of PE [30, 31, 32], those approaches in general lack robustness against process and measurement noise. In Figure 3, we show the application of the adaptive estimation approach [^4] Application of the adaptive observer from [32] requires to transform the system (3) to observable canonical form, consequently having a state of dimension $n=2$. We choose the parameters of the observer from [32] as $F=a\cdot I_{2}$ (with $a=0.99$ from the dynamics introduced in Section 3), $R=I_{2}$, and apply covariance resetting [32, Sec. 3.1] with $k_{0}=0.1$ and $k_{\min}=0.01$. from [32] to the academic example presented in Section 3, which is characterized by measurement noise and absence of PE. It is visible that the parameter estimate converges in the noise-free case (with $w=0$), resulting in small state estimation errors. In the noisy case, we observe a similar effect as for the MHE with prior on $\hat{\theta}_{t-M_{t}}$ as shown in Figure 2. In particular, the slowly diverging parameter estimates lead to large state estimation errors whenever there is enough excitation. In contrast, the proposed MHE approach is robustly stable (Definition 2) for any general nonlinear system whenever there exists any robustly stable estimator (Assumption 1). Hence, whenever it is in principle possible to design an adaptive observer with robust stability properties, then, Assumption 1 holds and the proposed MHE formulation yields robust stability properties. However, to the best of the author’s knowledge, existing adaptive observer designs do not have such properties for general nonlinear systems.
@@
The problem of parameter drift and the related bursting phenomena is well known in the adaptive control and estimation literature [64, 55, 53, 54] and the interplay between PE and robustness remains actively researched [36, 33]. We propose a simple constructive solution to this problem by introducing a regularization cost on the parameter estimate w.r.t. a constant initial parameter estimate $\bar{\theta}_{0}$. As shown in [65], such a regularization can be related to the classical $\sigma$-modification, compare [66, Chap. 8][67], which ensures boundedness/robustness irrespective of PE.
5.4 Functional Estimation
Functional estimation [68, 69, 50] is applied to estimate a function of the state $z_{t}=\phi(x_{t})$. In [50, Thm. 9], it was shown that robustly stable estimates of $z_{t}$ can be obtained using an FIE if the system admits a stable estimator (Assumption 1). The state estimation problem considered here is a special case of a functional estimation problem for the augmented system with state $\tilde{x}_{t}=\begin{bmatrix}x_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ with functional output $\tilde{z}_{t}=x_{t}$, compare [50, Rem. 1]. Since this functional estimation problem is nonlinear (even for linear systems), classical functional estimation approaches [68, 69] are not applicable. Corollary 4.3 recovers the FIE results for functional estimation from [50, Thm. 9]. In addition, such an FIE can even ensure (asymptotic) robust stability under a weaker asymptotic detectability condition [50, Thm. 9]. However, in contrast to the proposed MHE formulation with fixed horizon length, solving FIE is in general not computationally tractable.
The MHE design and theoretical analysis presented here show that it is possible to establish (practical) robust stability for functional estimation using MHE, but under certain (potentially restrictive) conditions. In particular, to establish (practical) robust stability of the estimates of the system state $x_{t}$, it is crucial that the (non-detectable) parameter $\theta$ is constant. Finding general conditions on the functional estimation problem which enable (practical) robust stability of MHE for functional estimation is an interesting topic of future research.
@@ 5.5 Verification of Assumption 1 in Practical Applications
with state $\begin{bmatrix}p_{t}^{\top}&v_{t}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{2\cdot n_{\mathrm{p}}}$, control input $u_{t}\in\mathbb{R}^{m}$, parameter $\theta\in\mathbb{R}^{n_{\theta}}$, measurement $y_{t}\in\mathbb{R}^{n_{\mathrm{p}}}$, process and measurement noise $w_{t}\in\mathbb{R}^{3\cdot n_{\mathrm{p}}}$, sampling time $\Delta t$, and $A^{-1}(p)$ Lipschitz continuous. Such a model (29) can describe the behavior of many mechanical systems where the position $p$ can be measured, e.g., quadrotors [40, Sec. V.B], and cars [70], compare also the presented numerical example in Section 6 below.
In the noise-free case (i.e., with $w=0$), the state of system (29) can be exactly recovered given two past outputs (32), even without knowledge of the parameters. Hence, a robustly stable state estimator can be constructed, i.e., Assumption 1 is satisfied [^5] In the noise-free case, the states can be recovered as $p_{t}=y_{t}$, $v_{t}=\frac{1}{\Delta t}A^{-1}(y_{t})\left(y_{t+1}-y_{t}\right)$. Due to Lipschitz continuity of $A^{-1}$, the state estimation error is bounded proportional to the magnitude of the process and measurement noise, i.e., Equation (9) holds with arbitrarily small positive constants $\lambda_{1},\lambda_{2},\lambda_{3}>0$ and $\epsilon=0$. However, note that this estimator is not causal, which slightly deviates from Definition 1.. On the other hand, the augmented state $\begin{bmatrix}p_{t}^{\top}&v_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ of system (29) is in general not uniformly detectable, compare, e.g., the numerical example in Section 6 below.
@@ 6.1 State Estimation for Car
We consider a dynamic bicycle model with simplified Pacejka tire force model [71, 70]. The continuous-time system is governed by following differential equations
@@
Table 1: Parameters of miniature car from [70]
$l_{\mathrm{r}}$ $l_{\mathrm{f}}$ $m$ $I_{\mathrm{z}}$
$0.038$ $0.052$ $0.181$ $5.05\cdot 10^{-4}$
$8.5$ $1.45$ $1.0$ $5.2$ $1.5$ $0.65$
@@
where $\alpha_{\mathrm{f}}$ and $\alpha_{\mathrm{r}}$ are the front and rear slip angles, and $B_{\mathrm{f}},\ B_{\mathrm{r}},\ C_{\mathrm{f}},\ C_{\mathrm{r}},\ D_{\mathrm{f}}$, and $D_{\mathrm{r}}$ are the Pacejka tire model parameters. The control input is $u=\begin{bmatrix}\delta&F_{\mathrm{x}}\end{bmatrix}^{\top}\in\mathbb{R}^{2}$ and the overall state $x=\begin{bmatrix}x_{\mathrm{p}}&y_{\mathrm{p}}&\psi&v_{\mathrm{x}}&v_{\mathrm{y}}&\omega\end{bmatrix}^{\top}\in\mathbb{X}=\mathbb{R}^{6}$. The parameters of the considered system (6.1) are stated in Table 1 and represent a miniature race car from [70]. The Pacejka parameters $D_{\mathrm{f}}$ and $D_{\mathrm{r}}$ are considered to be unknown, i.e., $\theta=\begin{bmatrix}D_{\mathrm{f}}&D_{\mathrm{r}}\end{bmatrix}^{\top}\in\Theta=\mathbb{R}^{2}$, with given prior $\bar{\theta}_{0}=2\cdot\theta$. The system dynamics (6.1) are discretized using Euler forward with sampling time $\Delta t=10ms$, and the resulting discrete-time model is subject to uniformly distributed additive process noise $w_{\mathrm{x}}\in\mathbb{W}_{\mathrm{x}}=\{w_{\mathrm{x}}\in\mathbb{R}^{6}|\|w_{\mathrm{x}}\|_{\infty}\leq 0.01\}$.
@@
Note that the augmented state $\begin{bmatrix}x^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ is not uniformly detectable as the friction parameters cannot be estimated for low/zero slip $\alpha_{\mathrm{f}}$ and $\alpha_{\mathrm{r}}$. In particular, if the car is driving straight ($\omega=0$, $\delta=0$, $v_{\mathrm{y}}=0$), the slip angles (31a)-(31b) and tire forces (31c)-(31d) are zero, i.e., no information about $\theta$ can be inferred. Consequently, no robustly stable estimator for the augmented state exists. On the other hand, the system (6.1)-(32) is in the form of (29) and hence the arguments in Section 5.5 apply, i.e., the derived stability results are applicable.
@@ 6.2 Numerical Estimation Results
Figure 4: Miniature race car example (Section 6): The top three plots show the states $x_{\mathrm{p}},\ y_{\mathrm{p}},\ \psi$ with corresponding sensor measurements and estimates resulting from a standard MHE approach with prior $\hat{\theta}_{t-M_{t}}$, the MHE approach without adaptation and fixed parameter value $\bar{\theta}_{0}$, and our proposed MHE approach with prior $\bar{\theta}_{0}$. The bottom three subplots show the estimation error for $v_{\mathrm{x}},\ v_{\mathrm{y}},\ \omega$ resulting from all three MHE approaches (dotted line) and a moving average thereof over $40$ time steps (solid line).
@@
Figure 5: Miniature race car example (Section 6): The top three subplots show the normalized parameter estimation errors resulting from a standard MHE approach with prior $\hat{\theta}_{t-M_{t}}$, the MHE approach without adaptation and fixed parameter value $\bar{\theta}_{0}$, and our proposed MHE approach with prior $\bar{\theta}_{0}$ (dotted line) and a moving average thereof over $40$ time steps (solid line). The bottom two subplots show the tire forces $F_{\mathrm{r}}$ and $F_{\mathrm{f}}$ and the corresponding estimates thereof resulting from all three MHE approaches.
@@
In the following, we consider a scenario where the car is driving straight ($u_{t}=\begin{bmatrix}0&0.01\end{bmatrix}^{\top}$), takes a right turn between time $3s$ and $6s$ ($u_{t}=\begin{bmatrix}-0.03&0.01\end{bmatrix}^{\top}$), and drives straight again, see the top three subplots of Figure 4. We compare the following three approaches: (i) the proposed MHE approach with prior weighting w.r.t $\bar{\theta}_{0}$ (Section 4), (ii) a standard MHE approach for state and parameter estimation using updated prior $\hat{\theta}_{t-M_{t}}$ (Section 2.3), and (iii) an MHE approach for state estimation without parameter adaptation, i.e., the proposed MHE formulation but with fixed parameter estimate $\hat{\theta}_{t}=\bar{\theta}_{0}$. All three estimators are implemented in Python using CasADi [72] and the solver IPOPT [73]. The code will be made available online. We choose $Q=\mathrm{diag}(10^{4}\cdot I_{6},25\cdot I_{2},10^{4})$ and $R=\mathrm{diag}(25\cdot I_{2},10^{4})$, reflecting the bounds in $\mathbb{W}$. Furthermore, we choose $\eta=0.9$, $P_{\mathrm{x}}=P_{2,\mathrm{x}}=I_{6}$, $P_{\theta}=P_{2,\theta}=I_{2}$, a horizon $M=20$, and $\bar{w}_{t}=0$ and $\bar{y}_{t}=y_{t}$ for all $t\in\mathbb{I}_{\geq 0}$. The system is initialized with $x_{0}=\begin{bmatrix}0&0&0&4&0&0\end{bmatrix}^{\top}$, while the initial estimate is chosen as $\bar{x}_{0}=\begin{bmatrix}0&0&0&3.8&0&0\end{bmatrix}^{\top}$.
@@
Figure 4 shows the ground truth states and state estimation errors resulting from all three MHE approaches. In Figure 5, the corresponding estimation errors of the unknown parameters $D_{\mathrm{r}}$ and $D_{\mathrm{f}}$, and the estimates of the tire forces $F_{\mathrm{r}}$ and $F_{\mathrm{f}}$ are shown. The first three states ($x_{\mathrm{p}}$, $y_{\mathrm{p}}$, and $\psi$) are accurately estimated from the noisy measurements. The proposed MHE approach results in smaller estimation errors for the longitudinal, lateral, and angular velocities, $v_{\mathrm{x}}$, $v_{\mathrm{y}}$, $\omega$, respectively, sometimes over one order of magnitude compared to the two other MHE schemes. The MHE approach with prior $\hat{\theta}_{t-M_{t}}$ leads to drifting parameter estimates while the car is driving straight. When the car is turning, this yields bad estimates of the tire forces and consequently large state estimation errors. The MHE approach without adaptation leads to inaccurate estimates of the tire forces during the turn, which also results in increased state estimation erros. In contrast, the proposed MHE approach prevents parameter drift when driving straight, while the parameter estimates are improved during the turn and consequently lead to accurate estimates of the tire forces and reduced state estimation errors.

### v2

@@ 5 Discussion
In this section, we discuss the relation of the MHE approach presented in Section 4 to existing methods for state and parameter estimation from the literature.
@@ 5.1 Standard Moving Horizon Estimation
Assuming perfect knowledge of model parameters $\theta$, $\delta$-IOSS is a necessary and sufficient condition for robustly stable state estimation [44]. Furthermore, MHE is robustly stable assuming $\delta$-IOSS and a sufficiently long horizon [41, 63]. Consequently, a robustly stable MHE approach can be designed whenever a robustly stable state estimator exists.
MHE is often applied to estimate an augmented system state combining both states and parameters [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Similarly to the state estimation case, the existence of a robustly (uniformly) stable estimator for the augmented system state implies stability of MHE with sufficiently long horizon, compare [51]. However, existence of a stable estimator in this setting generally requires satisfaction of some PE condition (cf. Remark 2.2), which is in general not given. Lack of PE can lead to parameter drift and unstable state estimates as shown in Section 3. We overcome this issue by replacing the standard prior weighting on the parameter with a regularization with respect to a (constant) prior parameter estimate $\bar{\theta}_{0}$. Only assuming existence of a robustly stable estimator (Assumption 1), the proposed MHE approach is shown to result in practically stable estimates of the state irrespective of PE (Theorem 4.3), i.e., the proposed MHE “performs well” whenever it is possible to design any observer satisfying the desired stability property.
@@ 5.2 Adaptations of Prior Weighting in Absence of PE
More recently, an MHE approach for joint state and parameter estimation with online verification of PE has been introduced in [51]. Thereby, the prior parameter estimate is kept fixed whenever there is insufficient PE as also done in the MHE approach proposed in Section 4.1. The MHE approach in [51] offers the possibility to adapt the prior parameter estimate in case of sufficient PE. However, practical application of the online verification of PE relies on a stricter assumption [51, Ass. 12], which limits applicability to weakly nonlinear systems with small parameter errors. In contrast, in this paper we rely on a necessary and sufficient condition (Ass. 1) to ensure practically robustly stable state estimates, even in cases of arbitrarily large initial parameter estimation errors.
@@ 5.3 Adaptive Estimation
Classically, adaptive observers are applied for state estimation in case of parametric uncertainty. While there exist adaptive observers with guarantees in the absence of PE [31, 32, 33], those approaches in general lack robustness against process and measurement noise. In Figure 3, we show the application of the adaptive estimation approach [^7] Application of the adaptive observer from [33] requires to transform the system (10) to observable canonical form, consequently having a state of dimension $n_{\mathrm{x}}=2$. We choose the parameters of the observer from [33] as $F=a\cdot I_{2}$ (with $a=0.99$ from the dynamics introduced in Section 3), $R=10\cdot I_{2}$, and apply covariance resetting [33, Sec. 3.1] with $k_{0}=0.1$ and $k_{\min}=0.01$. from [33] to the academic example presented in Section 3, which is characterized by measurement noise and absence of PE. It is visible that the parameter estimate converges in the noise-free case (with $w=0$), resulting in small state estimation errors. In the noisy case, we observe a similar effect as for the MHE with prior on $\hat{\theta}_{t-M_{t}}$ as shown in Figure 2. In particular, the slowly diverging parameter estimates lead to large state estimation errors whenever there is enough excitation. In contrast, the proposed MHE approach is practically robustly stable (Definition 2) for any general nonlinear system whenever there exists any robustly stable estimator (Assumption 1). Hence, whenever it is in principle possible to design an adaptive observer with robust stability properties, then, Assumption 1 holds and the proposed MHE formulation yields practical robust stability properties. However, to the best of the author’s knowledge, existing adaptive observer designs do not have such properties for general nonlinear systems.
@@
The problem of parameter drift and the related bursting phenomena is well known in the adaptive control and estimation literature [71, 60, 56, 59] and the interplay between PE and robustness remains actively researched [37, 34]. We propose a simple constructive solution to this problem by introducing a regularization cost on the parameter estimate w.r.t. a constant initial parameter estimate $\bar{\theta}_{0}$. As shown in [72], such a regularization can be related to the classical $\sigma$-modification, compare [73, Chap. 8][74], which ensures boundedness/robustness irrespective of PE.
@@ 5.4 Verification of Assumption 1 in Practical Applications
with state $\begin{bmatrix}p_{t}^{\top}&v_{t}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{2\cdot n_{\mathrm{p}}}$, control input $u_{t}\in\mathbb{R}^{n_{\mathrm{u}}}$, parameter $\theta\in\mathbb{R}^{n_{\theta}}$, measurement $y_{t}\in\mathbb{R}^{n_{\mathrm{p}}}$, process and measurement noise $w_{t}\in\mathbb{R}^{3\cdot n_{\mathrm{p}}}$, sampling time $\Delta t$, and $A^{-1}(p)$ Lipschitz continuous. Such a model (17) can describe the behavior of many mechanical systems where the position $p$ can be measured, e.g., quadrotors [41, Sec. V.B], and cars [75], compare also the presented numerical example in Section 6 below.
In the noise-free case (i.e., with $w=0$), the state of system (17) can be exactly recovered given two past outputs (20), even without knowledge of the parameters. Hence, a robustly stable state estimator can be constructed, i.e., Assumption 1 is satisfied [^8] In the noise-free case, the states can be recovered as $p_{t}=y_{t}$, $v_{t}=\frac{1}{\Delta t}A^{-1}(y_{t})\left(y_{t+1}-y_{t}\right)$. Due to Lipschitz continuity of $A^{-1}$, the state estimation error is bounded proportionally to the magnitude of the process and measurement noise, i.e., Equation (4) holds with arbitrarily small positive constants $\lambda_{1},\lambda_{2},\lambda_{3}>0$ and $\epsilon=0$. However, note that this estimator is not causal, which slightly deviates from Definition 1.. On the other hand, the augmented state $\begin{bmatrix}p_{t}^{\top}&v_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ of system (17) is in general not uniformly detectable, compare, e.g., the numerical example in Section 6 below.
@@ 6.1 State Estimation for Car
We consider a dynamic bicycle model with simplified Pacejka tire force model [76, 75]. The continuous-time system is governed by the following differential equations
@@
where $\alpha_{\mathrm{f}}$ and $\alpha_{\mathrm{r}}$ are the front and rear slip angles, and $B_{\mathrm{f}},\ B_{\mathrm{r}},\ C_{\mathrm{f}},\ C_{\mathrm{r}},\ D_{\mathrm{f}}$, and $D_{\mathrm{r}}$ are the Pacejka tire model parameters. The control input is $u=\begin{bmatrix}\delta&F_{\mathrm{x}}\end{bmatrix}^{\top}\in\mathbb{R}^{2}$, where $\delta$ is the steer angle of the front wheels and $F_{\mathrm{x}}$ the longitudinal force generated by the motor, and the overall state is $x=\begin{bmatrix}x_{\mathrm{p}}&y_{\mathrm{p}}&\psi&v_{\mathrm{x}}&v_{\mathrm{y}}&\omega\end{bmatrix}^{\top}\in\mathbb{X}=\mathbb{R}^{6}$. The parameters of the considered system (18) are stated in Table 1 and represent a miniature race car from [75]. The Pacejka parameters $D_{\mathrm{f}}$ and $D_{\mathrm{r}}$ are considered to be unknown, i.e., $\theta=\begin{bmatrix}D_{\mathrm{f}}&D_{\mathrm{r}}\end{bmatrix}^{\top}\in\Theta=\mathbb{R}^{2}$, with given prior $\bar{\theta}_{0}=2\cdot\theta$. The system dynamics (18) are discretized using Euler forward with sampling time $\Delta t=10ms$, and the resulting discrete-time model is subject to uniformly distributed additive process noise $w_{\mathrm{x}}\in\mathbb{W}_{\mathrm{x}}=\{w_{\mathrm{x}}\in\mathbb{R}^{6}|\|w_{\mathrm{x}}\|_{\infty}\leq 0.01\}$.
Table 1: Parameters of miniature car from [75]
$l_{\mathrm{r}}$ $l_{\mathrm{f}}$ $m$ $I_{\mathrm{z}}$
$0.038$ $0.052$ $0.181$ $5.05\cdot 10^{-4}$
$8.5$ $1.45$ 1.0-1.0 $5.2$ $1.5$ 0.65-0.65
@@
Note that the augmented state $\begin{bmatrix}x^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ is not uniformly detectable as the friction parameters cannot be estimated for low/zero slip $\alpha_{\mathrm{f}}$ and $\alpha_{\mathrm{r}}$. In particular, if the car is driving straight ($\omega=0$, $\delta=0$, $v_{\mathrm{y}}=0$), the slip angles (19a)-(19b) and tire forces (19c)-(19d) are zero, i.e., no information about $\theta$ can be inferred. Consequently, no robustly stable estimator for the augmented state exists. On the other hand, the system (18)-(20) is in the form of (17) and hence the arguments in Section 5.4 apply, i.e., the derived stability results are applicable.
@@ 6.2 Numerical Estimation Results
Figure 4: Miniature race car example (Section 6): The top three plots show the states $x_{\mathrm{p}},\ y_{\mathrm{p}},\ \psi$ with corresponding sensor measurements and estimates resulting from all three considered MHE approaches and an EKF. The bottom three subplots show the estimation error for $v_{\mathrm{x}},\ v_{\mathrm{y}},\ \omega$ resulting from all three MHE approaches and the EKF (dotted line) and a moving average thereof over $40$ time steps (solid line).
@@
Figure 5: Miniature race car example (Section 6): The top two subplots show the normalized parameter estimation errors resulting from all three considered MHE approaches and an EKF (dotted line) and a moving average thereof over $40$ time steps (solid line). The bottom two subplots show the tire forces $F_{\mathrm{r}}$ and $F_{\mathrm{f}}$ and the corresponding estimates thereof resulting from all three MHE approaches and the EKF.
Figure 6: Miniature race car example (Section 6): Signal to noise ratio (SNR) of minimum singular value of excitation matrix $PE_{t}$ divided by the maximum singular value of the empirical process noise covariance $\Sigma_{t}^{\mathrm{w}}$, compare Remark $2.$
In the following, we consider a scenario where the car is driving straight ($u_{t}=\begin{bmatrix}0&0.01\end{bmatrix}^{\top}$), takes a right turn between time $3s$ and $6s$ ($u_{t}=\begin{bmatrix}-0.03&0.01\end{bmatrix}^{\top}$), and drives straight again, see the top three subplots of Figure 4. Note that in this scenario, the signal to noise ratio (cf. Remark 2.2) is small before and after the turn, while it is large during the turn, compare Figure 6.
We compare the following four approaches:
  1. (i)
    MHE-proposed: the proposed MHE approach with prior weighting w.r.t. $\bar{\theta}_{0}$ as defined in Section 4,
  2. (ii)
    MHE-joint: a standard MHE approach for joint state and parameter estimation using updated prior $\hat{\theta}_{t-M_{t}}$ as stated in Section 2.3,
  3. (iii)
    MHE-no-adaptation: an MHE approach for state estimation without parameter adaptation, i.e., the proposed MHE formulation but with fixed parameter estimate $\hat{\theta}_{t}=\bar{\theta}_{0}$,
  4. (iv)
    EKF: an EKF to estimate the extended state $\tilde{x}_{t}=\begin{bmatrix}x_{t}^{\top}&\theta_{t}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{8}$ [17].
@@
All three MHE approaches are implemented in Python using CasADi [57] and the solver IPOPT [58]. The code is available [^9] https://gitlab.ethz.ch/ics/parametric-mhe online. For the MHE approaches, we choose $Q=\mathrm{diag}(10^{4}\cdot I_{6},25\cdot I_{2},10^{4})$ and $R=\mathrm{diag}(25\cdot I_{2},10^{4})$, reflecting the bounds in $\mathbb{W}$. Furthermore, we choose $\eta=0.9$, $P_{\mathrm{x}}=P_{2,\mathrm{x}}=I_{6}$, $P_{\theta}=P_{2,\theta}=I_{2}$, a horizon $M=20$, and $\bar{w}_{t}=0$ and $\bar{y}_{t}=y_{t}$ for all $t\in\mathbb{I}_{\geq 0}$. For the EKF, we choose the covariance matrices as $Q_{\mathrm{EKF}}=\mathrm{diag}(10^{-4}\cdot I_{6},2\cdot I_{2})$, $R_{\mathrm{EKF}}=R^{-1}/4$, and initialize the estimation error covariance matrix with $P_{0}=I_{8}$. The system is initialized with $x_{0}=\begin{bmatrix}0&0&0&4&0&0\end{bmatrix}^{\top}$, while the initial estimate is chosen as $\bar{x}_{0}=\begin{bmatrix}0&0&0&3.8&0&0\end{bmatrix}^{\top}$.
@@
Figure 4 shows the ground truth states and state estimation errors resulting from all three MHE approaches as well as the EKF. In Figure 5, the corresponding estimation errors of the unknown parameters $D_{\mathrm{r}}$ and $D_{\mathrm{f}}$, and the estimates of the tire forces $F_{\mathrm{r}}$ and $F_{\mathrm{f}}$ are shown. Note that in the lower three subplots of Figure 4 and in Figure 5, the dotted lines indicate the values at each time step, while the solid lines indicate a moving average thereof over $40$ time steps. As seen in Figure 4, the first three states ($x_{\mathrm{p}}$, $y_{\mathrm{p}}$, and $\psi$) are accurately estimated from the noisy measurements by all three MHE approaches and the EKF. The MHE-proposed results in smaller estimation errors for the longitudinal, lateral, and angular velocities, $v_{\mathrm{x}}$, $v_{\mathrm{y}}$, $\omega$, respectively, sometimes over one order of magnitude compared to the two other MHE schemes. While the EKF results in improved estimates of $v_{\mathrm{x}}$ compared to all three MHE approaches, the estimation errors of $v_{\mathrm{y}}$ and $\omega$ are larger during the right turn compared to the MHE-proposed. As seen in Figure 5, the MHE-joint leads to drifting parameter estimates while the car is driving straight, even with the considered unbiased noise distribution. When the car is turning, this yields bad estimates of the tire forces and consequently large state estimation errors. Similarly to the MHE-joint, the EKF leads to drifting estimates of $D_{\mathrm{f}}$ while the car is driving straight. In addition, the estimates of $D_{\mathrm{r}}$ resulting from the EKF have a significantly larger error compared to all three MHE approaches. The MHE-no-adaptation leads to inaccurate estimates of the tire forces during the turn, which also results in increased state estimation errors. In contrast, the MHE-proposed prevents parameter drift when driving straight, while the parameter estimates are improved during the turn and consequently lead to accurate estimates of the tire forces and reduced state estimation errors. Note that the state estimation errors in $v_{\mathrm{y}}$ and $\omega$ resulting from the MHE-proposed increase during the turn compared to when the car is driving straight. This is due to the biased parameter estimates affecting the dynamics of the car in this phase. However, they remain below the estimation errors resulting from the MHE-joint for joint state and parameter estimation and the MHE-no-adaptation.
The median computational time [^10] The time was taken on a laptop with 12-core Intel i7 processor and 32 GB of memory. of the MHE-joint, MHE-no-adaptation, and MHE-proposed are $9.3ms$, $7.6ms$, and $8.7ms$, respectively, while the considered time step is $10ms$. For embedded applications, the computational time can be significantly reduced by relying on tailored compiled solvers, see, e.g., recent experimental applications of a comparable MHE on a miniature race car [77].